Status Report

ISS On-Orbit Status 29 Jun 2002

By SpaceRef Editor
June 29, 2002
Filed under , ,

All ISS systems continue to function nominally,
except as noted previously or below

Docking of Progress M-46/8P this morning was smooth and
, with on-time capture at the SM aft port at 2:23am EDT.

After engagement of docking latches, CDR Valery Korzun and FE-2
Sergei Treschev performed interface leak check, then opened hatches
at about 5:30am and installed the QD (quick disconnect) screw clamps.
 Korzun took air samples in the Progress with the Russian AK-1M
air sampler, deactivated the cargo ship’s systems, and installed the
air duct between SM and Progress.  Later, Korzun and Treschev
deconfigured the Ku-band video setup used for the docking.

For the crew, the rest of the weekend, after the arrival of the fully
loaded resupply ship, holds the excitement of a treasure hunt through
all their new supplies.

FE-1 Peggy Whitson reviewed procedures for checking out the new MSG
(Microgravity Science Glovebox) and later completed initial checkout,
inspection of mechanical parts and filter installation on the MSG.
The task included loosening of the launch bolts, but installation of
rack cabling will be done later this week as a task list item.

FE-2 Sergei Treschev was scheduled for transferring the hardware of
the MEPS (microencapsulation electrostatic
processing) payload from EXPRESS Rack 3 (ER3) to its operational
location in ER1.  Since this required swapping ofamickers and
contents, the ground needs to update the IMS (inventory management
system) database accordingly.

POC (Payload Operations Center) has added deactivation of
PCG-STES PCAM (Protein Crystal Growth-Single Locker Thermal Enclosure
System / Protein Crystallization Apparatus for Microgravity) cylinder
#3 to the task list.  [Based on ground testing, as well as
previous vapor equilibration studies, the science team has determined
that the crystals produced in this experiment will be best preserved
by early deactivation of this cylinder.]

Treschev terminated regeneration of absorbent bed #1 of the SM
BMP micropurification unit.

Whitson performed the daily routine payload status check, while
Treschev did the SOSh life support systems maintenance and Korzun the
regular inspection of the BRPK-2 water condensate separator unit.

All crewmembers performed their physical exercise.

Science Update (Expedition Five — 3rd):
Increment lead scientist Dr. Vic Cooley reported that the
space-grown ADVASC soybeans "look very healthy and robust".
 He thanked Sergei and Peggy for troubleshooting the ZCG backup
hard-drive and getting the payload going: "ZCG and PCG-STES
crystals should be growing bigger by the hour". ARIS performed
excellently isolating the ZCG payload through the critical first 20

Extra-Vehicular Activity Radiation Monitors (EVARM):  The
UF-2 EVA data along with the background radiation readings completed
6/28 have been received by the PI and are currently being

GASMAP/Pulmonary Function in Flight (PuFF):  First
inflight session has been completed.  Ku coverage allowed the
ground to see about half of the data.  What they saw looked good
and the live video was "truly impressive".  The PuFF
team sent thanks for the crew’s hard work and looks forward to a
productive increment.

Renal (Kidney) Stone Experiment:   First data
collection session is scheduled for Week 5. This fulfills the 30 day
requirement for the first collection.

Interactions (NTXN):  A big ‘thank you’ from the
Interactions team to the crewmembers for performing their session
past week.

Human Research Facility/Workstation (HRF WS):

Commercial refrigerator incubator module-commercial Stelsys
Experiment is complete and samples are stowed in

Microencapsulation Electrostatic
Processing (MEPS):
 Hardware transfer from Express
Rack 3 to Express Rack 1 took place today.  Hardware activation
will most likely occur during Week 6.

Advanced Astroculture (ADVASC):  The ADVASC hardware
continues to operate normally.  Chamber parameters are precisely
controlled at the predefined conditions.  The soybeans are
healthy and growing well.  They are currently at the V1 stage
with mature unifoliolate leaves and developing trifoliolate leaves
(first set). The plants are growing on the expected schedule.

Solidification Using a Baffle in Sealed Ampoules (SUBSA):
ctivities for SUBSA are scheduled next week (Week 4) and
operations in Week 5.

Pore Formation and Mobility Investigation (PFMI):  New
payload for Increment 5 has arrived. Will be performed during Stage

Space Acceleration Measurement System (SAMS):  The ground
is investigating a new 60Hz signal seen in spectrogram from head
121f05.  Preliminary investigations indicate HRF/PuFF/GASMAP
activities.  Today’s recordings included Progress 8P docking and
MSG checkout.

Microgravity Acceleration Measurement System (MAMS):
 MAMS captured progress 7P undocking and attitude transitions.
 HiRAP data downlink was enabled for 8P Progress docking.

Protein Crystal Growth-Single Locker Thermal Enclosure System
 The STES-8 unit is running nominal at 22.0 deg
C since transfer.

Materials ISS Experiment (MISSE):  In progress. Deployed
outside. Nominal and collecting data.

Educational Payload Operations (EPO):  Planned. Will be
performed during Stage 9A.

EarthKAM:   Planned.  Will be performed during
Stage 9A.  
All EarthKAM images are available for public access on the EarthKAM
data system at:

Zeolite Crystal Growth (ZCG):  After troubleshooting
6/24-26, ZCG still could not get the back-up computer to boot.
 The ground activated on 6/27 at 10:38am EDT using only the
primary computer.   The PI has determined that this
particular set of samples will only run for 15 days instead of the 25
that were originally planned. The hardware is currently nominal
except for the backup CPU (central processor unit).

ARCTIC Refrigerator/Freezer 1 (ARCTIC-1):  ARCTIC-1 is
performing nominally at the -25.0 C set-point and the STELSYS samples
are kept frozen.

ARCTIC Refrigerator/Freezer 2 (ARCTIC-2):  The ARCTIC-2
unit is currently set at +4.0 C.  Since STELSYS activities are
now complete and there are no other payloads that require sample
stowage at this temp.  The unit is not needed for the remainder
of Increment 5; however, the ARCTIC team is investigating a possible
water flow problem.

Commercial Generic Bioprocessing Apparatus (CGBA):  Will
be performed during Stage 9A.

Crew Earth Observation (CEO):   Today’s target areas
were E. Mediterranean Dust and Smog (on this pass the crew
was to wait until they reached the lower Balkan Peninsula and then
look obliquely to the left of track for heavy smog reported over the
northern Aegean Sea and eastern Turkey, continuing to look eastward
over the Med to near the coast of Israel)
, Gulf of St.
(a major storm system has just moved through this
region. Of interest: looking obliquely to the left of track for fresh
snow cover inland and sea ice fragments offshore from Labrador)
Lakes of the Eastern Sierras (although lighting was low,
this pass had good weather and offered near-nadir views of both Mono
and Walker Lakes.  Crew to use the long lenses of the ESC
[electronic still camera] to detect shoreline changes of these water
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 11:54 am

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (16-amp mode).
Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum
pump failed).  U.S. CDRA CO2 scrubber is Off (failed).  BMP
Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in
Purify mode.

SM Working Compartment:  Pressure (mmHg) — 756,
temperature (deg C) — 26.3, ppO2 (mmHg) — 137.1, ppCO2 (mmHg) —
SM Transfer Compartment:  Pressure (mmHg) — 762,
temperature (deg C) — 20.3.
FGB Cabin:  Pressure (mmHg) — 752, temperature (deg C)
— 20.0.
Node:  Pressure (mmHg) — 748.34, temperature (deg C) —
24.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 755.67, temperature (deg C)
— 25.1, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 750.57,
temperature (deg C) — 23.9; shell heater temp (deg C) — 24.4, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 24.8
PMA-2:  Shell heater temp (deg C) — 16.9

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). New MSA (mass spectrometer
assembly) and VGA (verification gas assembly) have been installed,
but not activated yet.

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B in
Autotrack mode.
SM batteries: Battery #1 reported failed (6/21). All other batteries
(7) are in "Partial charge" mode.
FGB: Battery #6 is offline (ROM mode).  All other batteries (4)
are in ?Partial charge? mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Attitude Source:
State vector — US GPS (SIGI string 1)
Attitude — Russian
Angular rates — US RGA1 (rate gyro assembly 1)

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is powered Off (due
to consistently reaching upper operational temperature during

SSRMS/Canadarm2 at MBS PDGF (mobile base system/power &
data grapple fixture), with Keep Alive power on both strings.
RWS (robotics workstations): Cupola is Active, Lab RWS is Off.

SpaceRef staff editor.